In his essay “Don’t Blame the Eater,” David Zinczenko says he sympathizes with a group of children who are suing McDonald’s for making them fat. Drawing on his own experiences as a “1980’s latchkey kid,” he describes how easy it is for teenagers to put on weight with a steady diet of fast food meals. Part of the problem, he argues, is that nutritional information about fast food is often either unavailable or hard to interpret. For instance, at many fast food restaurants, when added ingredients and serving sizes are taken into account, even seemingly “healthy” menu options such as salads can include a shockingly large number of calories. In his opinion, the sheer number of fast food restaurants is also part of the problem, since if you “[d]rive down any throughfare in America, ... you’ll see one of our country’s more than 13,000 McDonald’s restaurants. Now drive back up the block and try to find someplace to buy a grapefruit.” In sum, Zinczenko argues that, especially for teenagers, unhealthy and fattening fast food is almost unavoidable.

In his essay “Don’t Blame the Eater,” David Zinczenko claims that the prevalence of fast food and the lack of healthier food alternatives is causing an epidemic of teenage obesity in present-day America. This is a serious issue, he argues, because of the consequences for society of increased rates of incidence of Type 2 diabetes, a disease closely associated with obesity. Drawing on his own experiences as a “1980’s latchkey kid,” he describes how easy it is for teenagers to put on weight with a steady diet of fast food meals, claiming that “by age 15, I had packed 212 pounds of torpid teenage tallow onto my once lanky 5-foot-10 frame.” Although he does not say so directly, Zinczenko seems to be implying here that there are hard and fast limits to the weight that a person of a certain height can carry safely. This raises a question, though. Is being “overweight” a medical condition that can be precisely defined, or is it at least partly a matter of perception and possibly dependent on cultural factors?